Operations Report – March 13th

Crew 190 Operations Report 13th March 2018
SOL: 2
Name of person filing report: Bastien BAIX
Non-nominal systems: –
Notes on non-nominal systems: –

Generator (hours run): Turned off at 10:36 am and turned on at 7:30 pm.
Solar – SOC 84% (Before generator is run at night)
Diesel – ~45%
Propane – 82%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.
Water (static) – ~450 gallons
Water (third tank) – 500 gallons.
Trailer to Static Pump used – NO
Water (loft) – Static to Loft Pump used – YES
Water Meter: 132 074.2
Toilet tank emptied: NO

ATV’s Used: 2 & 3
Oil Added: NO
ATV Fuel Used: NO
Hours the ATVs were Used today: 1 hour
Notes on ATVs: –

Deimos rover used: NO
Hours: 116.6
Beginning charge: 100%
Ending charge: –
Currently charging: NO
Notes on Deimos rover: flat tyre

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 24.8
Beginning charge: 100%
Ending charge: –
Currently charging: NO

Opportunity rover used: NO
Hours: 20.6
Beginning charge: 100%
Ending charge: –
Currently charging: YES

Curiosity rover used: YES
Hours: 17,2
Beginning charge: 76%
Ending charge: 100%
Currently charging: YES

HabCar used and why, where: –
General notes and comments: –
Summary of internet: ~ 200 Mb remaining
Summary of suits and radios: –
Summary of Hab operations: Ariane & Bastien working on their experiments
Summary of GreenHab operations: Mario & Fred working on their experiments. Aquaponic fish seedlings did not sprout. Suspected due to cold water. Pump turned on and hydroponic solution added to rectify.
Summary of ScienceDome operations: Martin, Sophie, Maximilien &
Michael working on their experiments.
Summary of RAMM operations: –
Summary of health and safety issues: HSO first report submitted

Questions, concerns and requests to Mission Support: –

Copyright © The Mars Society. All rights reserved. | Main Site